OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos x = sin x

OpenStudy (anonymous):

Sketch the graph.

OpenStudy (anonymous):

\[\frac{ \pi }{ 4 }..and..\frac{ 5\pi }{ 4 }\]

OpenStudy (anonymous):